OpenStudy (anonymous):

j(x)=ln(2x+1)e^(3x) Determine the derivative function. Please help! I do not know where to begin.

OpenStudy (mathmale):

Start by identifying what type function you have here. I see immediately that it's a product: the product of ln(2x+1) and e^(3x). So, begin finding the derivative by applying the PRODUCT RULE.

OpenStudy (johnweldon1993):

And the product rule states \[\large f(x)g'(x) + f'(x)g(x)\] Let f(x) = ln(2x + 1) and g(x) = e^(3x)

OpenStudy (anonymous):

okay, f'(x)= [3e^(3x) \ln(2x+1) +(2e^(3x))/(2x+1)] Does this look right?

OpenStudy (johnweldon1993):

Looks correct to me!